1. Write the given differential equation in the form L(y) = g(x), where L is a linear differential operator with constant coefficients. If possible, factor L. (Use D for the differential operator.)

y''' + 14y'' + 49y' = ex

_______ y= e^x

2. Find a linear differential operator that annihilates the given function. (Use D for the differential operator.)

1 + 3x − 4x3
1
Expert's answer
2020-06-30T17:57:03-0400

Solution:

  1. If Dy:=y' then D2y=y'',D3y=y''' and y''' + 14y'' + 49y' =D3y+14D2y+49Dy

We have equation D3y+14D2y+49Dy=ex

and (D3+14D2+49D)y=ex.

Then we can write the given equation in the form

L(y)=ex with operator L=D3+14D2+49D

After factorization we obtain

D3+14D2+49D=D(D2+14D+49)=D(D+7)2

Answer:D(D+7)2=ex

2.1+3x-4x3= (1+3x-4x3)e0x

Polynom 1+3x-4x3 has degree 3.

Then "\lambda" =0, k=4

Construct polynom which has root λ=0 of multiplicity 4:

λ4=0

Аnd appropriate linear differential operator with constant coefficients is L=D4

Answer:D4